OpenStudy (anonymous):

Solve the system of equations: what is the x-value for the solution? 2x + 7y = 3 and x = 1 - 4y A. -1 B. 5 C. 4 D. 3

OpenStudy (ricardovg):

7y + 2x= 3 4y + x = 1 --------------- 7y + 2x= 3 2(4y + x) = 2 => 8y + 2x = 2 ---------------- subtracting y + 0 = -1 => y=-1 and x=5